I am having a problems with these two problems and wondering if there was someone that could guide me on how to solve this problem. Thank you
Solve the following equation for X.
2/5=X-2/20


Answer by stanbon(75887)

Solve the following equation for X.
2/5=X-2/20
Multiply both sides by 20 to get of the denominators:
8 = 20x - 2
20x = 10
x=1/2
================
Cheers,
Stan H.